The figures below show the graphs of the exponential functions f(a) and g(x)\mid , and the linear function, h(x)\mid . The functionf(x)\mid has y-intercept 0.75 and goes through the point (1,3). The function g(x)\mid has y-intercept 2 and goes through the point (2,2/9). The functionh(x)\mid has y-intercept 5 and goes through the point (a,a+5).

(a) Find a formula for f(x)=\mid

(b) Find a formula for g(x)=\mid

(c) Find a formula for h(x)=\mid

(d) Find the exact value(s) of x=\mid such that f (x) f (x) = g (x)\mid

If there is more than one solution, enter your answers as a comma separated list.

x=\mid
Answer #1

For the first two, we use f(x) = a bx

For the first graph , points are (0, 0.75) , (1,3)

That gives

3= 0.75(b)1

b= 3/ 0.75 = 4

Equation is y= 0.75(4)x

For the second graph, points are (0,2),(2, 2/9)

2/9 = 2(b)2

1/ 9 = b2

b= 1/3

Equation is y = 2(1/3)x

For the third equation , points are (0,5),(a,a+5)

Slope = (y2-y1)/(x2-x1) = (a+5-5)/(a-0)=1

Equation of line is y= x+5

Part d. f(x)= g(x)

0.75(4)x=2(1/3)x

(4/1/3)x=2/0.75

12x=8/3

x log 12 = log(8/3)

x =log(8/3)/log(12)